# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 3.1 Inner region near $x=x_{0}$
$\xi\equiv\frac{x-x_{0}}{(3t)^{1/3}}=\mathcal{O}(1).$ (23)
Using this new similarity-type variable, we can rewrite (9) as
$u_{0}\sim\frac{A_{0}}{(3t)^{2/3}\xi^{2}},\quad u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(3t)^{5/3}\xi^{5}},$ (24)
@@
$u=\frac{1}{(3t)^{2/3}}f(\xi,t).$ (25)
By substituting (25) into (1), we find that $f$ satisfies the pde
$3tf_{t}-2f-\xi f_{\xi}+6ff_{\xi}+f_{\xi\xi\xi}=0.$ (26)
To leading order we write $f\sim f_{0}(\xi)$, so that (26) and (24) combine to give our inner problem
$-2f_{0}-\xi f_{0}^{\prime}+6f_{0}f_{0}^{\prime}+f_{0}^{{}^{\prime\prime\prime}}=0,$ (27)
$f_{0}\sim\frac{A_{0}}{\xi^{2}}+\frac{4A_{0}(A_{0}+2)}{\xi^{5}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty.$ (28)
Note the direction of the far-field condition (28) comes from matching back down towards the real axis. We expect (28) to apply in a broader sector of the $\xi$ plane, as determined by our subsequent analysis. Further, the extent to which (28) may need to be supported by other far-field conditions will be explored later.
We comment here that the similarity reduction (25) and the corresponding differential equations (26) and (27) are also used to derive the large-time asymptotics for the KdV problem (1) [1]. In that case, the similarity variable is $\xi\equiv x/(3t)^{1/3}$ (instead of (23)), and the appropriate solution of (27) is valid on the real line for $x=\mathcal{O}(t^{1/3})$ as $t\rightarrow\infty$. The relevant boundary conditions for (27) come from matching to an outer region as $\xi\rightarrow+\infty$. All of these variables for the large-time asymptotics are real valued. In contrast, our study of (25)-(27) is relevant for the region of the complex plane $x-x_{0}=\mathcal{O}(t^{1/3})$ as $t\rightarrow 0^{+}$. The boundary condition (28) is completely different to the matching condition used for the large-time asymptotics. The variables $f$ and $\xi$ in our study of (25)-(27) are complex valued.
The third-order nonlinear ordinary differential equation (ode) problem (27)–(28) is difficult to analyse in this form, but much progress can be made by converting (27) to PII in the usual way [52] (see the following subsection). Ultimately, one of the goals of this exercise is to determine the singularities of $f_{0}(\xi)$ in the $\xi$ plane. Our analysis then predicts that for any given singularity of $f_{0}$, which we call $\xi_{0}$, there is a double pole of $u(x,t)$ at $x=s(t)$ that emerges from $x_{0}$ like
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (29)
Before we proceed, there is an immediate exact solution for $A_{0}=-2$, which is
$f_{0}=-\frac{2}{\xi^{2}},\quad A_{0}=-2.$ (30)
@@ 3.2 Reduction to Painlevé II
$f_{0}=F^{\prime}-F^{2},$ (31)
we find that (27) is transformed to
$(F^{\prime\prime}-\xi F-2F^{3})^{\prime\prime}-2F(F^{\prime\prime}-\xi F-2F^{3})^{\prime}=0.$ (32)
Before we consider (32) further, note that given the far-field condition (28) and the change of variable (31), we must have
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (33)
@@
$\alpha={\textstyle\frac{1}{2}}(-1\pm\sqrt{1-4A_{0}}).$ (34)
@@
Returning to (32), clearly one option is
$F^{\prime\prime}-\xi F-2F^{3}=\,\mbox{constant},$ (35)
which is PII. To determine the constant, we apply the far-field condition (33), which implies it is $\alpha$. We are free to proceed with either sign in (34); each choice will involve different solutions for $F$ but the same result for $f_{0}$ via (31). We choose the plus sign. The other option, namely $F^{\prime\prime}-\xi F-2F^{3}\neq\,\mbox{constant}$, is not applicable as it is inconsistent with (33).
In summary, we conclude that (27)–(28) reduces to
$F^{\prime\prime}=2F^{3}+\xi F-\alpha,$ (36)
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (37)
@@
$\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{1-4A_{0}}).$ (38)
As mentioned above, (36) is PII with the constant $\alpha$ related to our key parameter $A_{0}$ (sometimes (36) is referred to as the inhomogeneous PII equation, with the homogenous version arising from $\alpha=0$). Motivated by the examples discussed in the Introduction, we are mostly focused on real and negative values of $A_{0}$, which correspond to $\alpha>0$. For example, the $N$-soliton solutions with $A_{0}=-N(N+1)$ correspond to $\alpha=N$ (via the rational solutions summarised in subsection 3.4). For $0\leq A_{0}\leq 1/4$, we have $1/2\leq\alpha\leq 0$, and for $A_{0}>1/4$ we have complex values of $\alpha$. These parameter intervals are also of some interest, although we shall restrict out exploration to real values of $\alpha$. Finally, without further investigation (or a background in studying PII) it is not obvious whether or not (37) is sufficient to determine our solution uniquely, which is an issue we explore in the next section.
@@ 3.3 Liouville–Green (WKB) analysis and Stokes phenomenon
For simplicity, we assume for the moment that the initial condition $u_{0}(x)$ has a single double pole in the upper half plane at $x=x_{0}$ (like (6) does). We wish to study (36) subject to (37). The solution we are after will satisfy this far-field condition in a broader sector that includes the negative imaginary direction. To determine this sector, and also to determine whether (37) is sufficient to specify our solution uniquely, we shall linearise about (37).
@@
which is related to the modified Bessel’s equation via a change of variables. The two possible linearly independent solutions have the far-field behaviour $J\sim\mathrm{const}\,\xi^{-1/4}\,\mathrm{e}^{\pm 2\xi^{3/2}/3}$. Therefore we have
@@
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty.$ (39)
This exercise demonstrates that (37) is acting as one boundary condition only and to specify the solution of (36)–(37) uniquely we need to fix $\sigma_{1}$ in (39), which is hidden beyond all orders of the algebraic expansion.
At this stage, we emphasise that (36) with (39) combine to give a one-parameter family of solutions of the type studied at length by Boutroux [14] and by others, including in [38, 40, 51, 53, 54]. We note that we are able to write out the full algebraic series in (39) as
$F\sim\frac{\alpha}{\xi}\sum_{n=0}^{\infty}\frac{b_{n}}{\xi^{3n}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (40)
@@
$b_{n+1}=(3n+1)(3n+2)b_{n}-2\alpha^{2}\sum_{\begin{subarray}{c}i,j,k=0\\ i+j+k=n\end{subarray}}^{n}b_{i}b_{j}b_{k},\quad b_{0}=1.$ (41)
@@
$b_{n}\sim\frac{\sin\pi\alpha}{\alpha\pi^{3/2}}\left(\frac{3}{2}\right)^{2n+1/2}\Gamma\left(2n+{\textstyle\frac{1}{2}}\right)\quad\mbox{as}\quad n\rightarrow\infty$ (42)
(see also [38]). A key point is that for $\alpha\in\mathbb{N}$, the series (40) is convergent (see section 3.4 below), while otherwise it is divergent.
To proceed with our argument about the one-parameter family of solutions to (36) with (39), we explore broader far-field sectors in the $\xi$ plane, referring to figure 7. With $\xi=\rho\mathrm{e}^{\mathrm{i}\theta}$, then let the direction $\xi\rightarrow-\mathrm{i}\infty$ be $\theta=-\pi/2$. For any $\sigma_{1}\neq 0$, then we would have $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ decaying along the ray $\theta=-\pi/2$. Along a path $\rho=\,\mathrm{const}$ (with $\rho\gg 1$), as $\theta$ increases from $\theta=-\pi/2$ this term would increase in size until it was $\mathcal{O}(1)$ at the anti-Stokes line $\theta=-\pi/3$ (dashed line in figure 7). For $\theta>-\pi/3$, the term would be exponentially growing.
@@
In order to eliminate the inclusion of an exponentially growing term in the sector $\pi<\theta<0$, we need to force $\sigma_{1}=0$ in (39). In that case, both exponentials $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ and $\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$ are absent along $\theta=-\pi/2$. As $\theta$ increases from $\theta=-\pi/2$, we see that $\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$ is switched on at the Stokes line $\theta=0$ (solid blue line) and becomes of $\mathcal{O}(1)$ at $\theta=\pi/3$. Transversing the other way, as $\theta$ decreases from $\theta=-\pi/2$, the exponential $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ is switched on at the Stokes line $\theta=-2\pi/3$ (solid blue line) and becomes of $\mathcal{O}(1)$ at $\theta=-\pi$. Putting it together, a complete description of our inner problem is
$F^{\prime\prime}=2F^{3}+\xi F-\alpha,$ (43)
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad|\xi|\rightarrow\infty,\quad-2\pi/3<\mathrm{arg}(\xi)<0$ (44)
$\sigma_{1}=0,\quad-2\pi/3<\mathrm{arg}(\xi)<0,$ (45)
where it is understood that the ellipsis represent further terms in the divergent power series (40); that is, there is a one-parameter family of solutions that satisfies (43)–(44), but the specific solution we are after is selected by enforcing (45). Furthermore, the unique solution to (43)–(45) should be pole-free region in the far field of the sector $\pi<\theta<\pi/3$ (bounded by active anti-Stokes lines at $\theta=-\pi$ and $\theta=\pi/3$) and the power-series part of the far-field condition, namely
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad|\xi|\rightarrow\infty,$ (46)
@@
Finally, from a practical (numerical) perspective, perhaps the best way to enforce (44)–(45) (which count as two boundary conditions) is to apply the far-field condition (46) along the ray $\theta=-\pi/3$ (the inactive anti-Stokes line). The reason is that for any $\sigma_{1}\neq 0$, the condition (46) would not apply along $\theta=-\pi/3$ (since, as already explained, the term $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ would be $\mathcal{O}(1)$ along this ray). Therefore, applying (46) along $\theta=-\pi/3$ has the effect of killing off $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ (as well as the other exponential $\sigma_{2}\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$). As an example of a numerical solution of (43)–(45) that is found by enforcing (46) along $\theta=-\pi/3$, we show in figure 8 an image computed for the case $\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{5})$ using the algorithm outlined by Fornberg & Weideman [40]. Here the yellow dots indicate poles of the solution, which are aligned in arrays. The red lines indicate anti-Stokes lines. Note that for this value of $\alpha$, there are no poles in the third or fourth quadrant. We discuss other numerical solutions of (43)–(45) below in subsection 3.5.
@@
Figure 8: An image showing $|F|$ for a numerical solution of Painlevé II with $\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{5})$ using the algorithm from Fornberg & Weideman [40]. The “initial conditions” obtained numerically using this scheme are $F(0)\approx 0.5941+1.0289\mathrm{i}$ and $F^{\prime}(0)\approx 0.7995-1.3848\mathrm{i}$. The yellow dots represent the pole field for this solution. The two active anti-Stokes lines are shown in red.

### v2

@@ 3.1 Inner region near $x=x_{0}$
$\xi\equiv\frac{x-x_{0}}{(3t)^{1/3}}=\mathcal{O}(1).$ (3.1)
Using this new similarity-type variable, we can rewrite (2.3) as
$u_{0}\sim\frac{A_{0}}{(3t)^{2/3}\xi^{2}},\quad u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(3t)^{5/3}\xi^{5}},$ (3.2)
@@
$u=\frac{1}{(3t)^{2/3}}f(\xi,t).$ (3.3)
By substituting (3.3) into (1.1), we find that $f$ satisfies the pde
$3tf_{t}-2f-\xi f_{\xi}+6ff_{\xi}+f_{\xi\xi\xi}=0.$ (3.4)
To leading order we write $f\sim f_{0}(\xi)$, so that (3.4) and (3.2) combine to give our inner problem
$-2f_{0}-\xi\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}+6f_{0}\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}+\frac{\mathrm{d}^{3}f_{0}}{\mathrm{d}\xi^{3}}=0,$ (3.5)
$f_{0}\sim\frac{A_{0}}{\xi^{2}}+\frac{4A_{0}(A_{0}+2)}{\xi^{5}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty.$ (3.6)
Note the direction of the far-field condition (3.6) comes from matching back down towards the real axis. We expect (3.6) to apply in a broader sector of the $\xi$ plane, as determined by our subsequent analysis. Further, the extent to which (3.6) may need to be supported by other far-field conditions will be explored later.
We comment here that the similarity reduction (3.3) and the corresponding differential equations (3.4) and (3.5) are also used to derive the large-time asymptotics for the KdV problem (1.1) [3]. In that case, the similarity variable is $\xi\equiv x/(3t)^{1/3}$ (instead of (3.1)), and the appropriate solution of (3.5) is valid on the real line for $x=\mathcal{O}(t^{1/3})$ as $t\rightarrow\infty$. The relevant boundary conditions for (3.5) come from matching to an outer region as $\xi\rightarrow+\infty$. All of these variables for the large-time asymptotics are real valued. In contrast, our study of (3.3)-(3.5) is relevant for the region of the complex plane $x-x_{0}=\mathcal{O}(t^{1/3})$ as $t\rightarrow 0^{+}$. The boundary condition (3.6) is completely different to the matching condition used for the large-time asymptotics. The variables $f$ and $\xi$ in our study of (3.3)-(3.5) are complex valued.
The third-order nonlinear ordinary differential equation (ode) problem (3.5)–(3.6) is difficult to analyse in this form, but much progress can be made by converting (3.5) to PII in the usual way [59] (see the following subsection). Ultimately, one of the goals of this exercise is to determine the singularities of $f_{0}(\xi)$ in the $\xi$ plane. Our analysis then predicts that for any given singularity of $f_{0}$, which we call $\xi_{0}$, there is a double pole of $u(x,t)$ at $x=s(t)$ that emerges from $x_{0}$ like
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (3.7)
Before we proceed, there is a simple exact solution of (3.5)–(3.6), namely
$f_{0}=-\frac{2}{\xi^{2}},\quad A_{0}=-2.$ (3.8)
@@ 3.2 Reduction to Painlevé II
$f_{0}=\frac{\mathrm{d}F}{\mathrm{d}\xi}-F^{2},$ (3.9)
we find that (3.5) is transformed to
$\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}\left(\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}-\xi F-2F^{3}\right)-2F\frac{\mathrm{d}}{\mathrm{d}\xi}\left(\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}-\xi F-2F^{3}\right)=0.$ (3.10)
Before we consider (3.10) further, note that given the far-field condition (3.6) and the change of variable (3.9), we must have
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (3.11)
@@
$\alpha={\textstyle\frac{1}{2}}(-1\pm\sqrt{1-4A_{0}}).$ (3.12)
@@
Returning to (3.10), clearly one option is
$\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}-\xi F-2F^{3}=\,\mbox{constant},$ (3.13)
which is PII. To determine the constant, we apply the far-field condition (3.11), which implies it is $\alpha$. We are free to proceed with either sign in (3.12); each choice will involve different solutions for $F$ but the same result for $f_{0}$ via (3.9). We choose the plus sign. The other option, namely $F^{\prime\prime}-\xi F-2F^{3}\neq\,\mbox{constant}$, is not applicable as it is inconsistent with (3.11).
In summary, we conclude that (3.5)–(3.6) reduces to
$\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}=2F^{3}+\xi F-\alpha,$ (3.14)
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (3.15)
@@
$\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{1-4A_{0}}).$ (3.16)
As mentioned above, (3.14) is PII with the constant $\alpha$ related to our key parameter $A_{0}$ (sometimes (3.14) is referred to as the inhomogeneous PII equation, with the homogenous version arising from $\alpha=0$). Motivated by the examples discussed in the Introduction, we are mostly focused on real and negative values of $A_{0}$, which correspond to $\alpha>0$. For example, the $N$-soliton solutions with $A_{0}=-N(N+1)$ correspond to $\alpha=N$ (via the rational solutions summarised in subsection 3.5). For $0\leq A_{0}\leq 1/4$, we have $1/2\leq\alpha\leq 0$, and for $A_{0}>1/4$ we have complex values of $\alpha$. These parameter intervals are also of some interest, although we shall restrict our exploration to real values of $\alpha$. Finally, without further investigation (or a background in studying PII) it is not obvious whether or not (3.15) is sufficient to determine our solution uniquely, which is an issue we explore in subsection 3.4.
3.3 Reduction to Painlevé 34
In our work, we have applied the Miura-transformation (3.9) to (3.5)–(3.6) so that our inner problem is described by the PII problem (3.14)–(3.16). Alternatively, we can set
$f_{0}=g_{0}+{\textstyle\frac{1}{2}}\xi,$ (3.17)
so that
$g_{0}+2\xi\frac{\mathrm{d}g_{0}}{\mathrm{d}\xi}+6g_{0}\frac{\mathrm{d}g_{0}}{\mathrm{d}\xi}+\frac{\mathrm{d}^{3}g_{0}}{\mathrm{d}\xi^{3}}=0.$
Multiplying by $g_{0}$ and integrating, we find
$g_{0}\frac{\mathrm{d}^{2}g_{0}}{\mathrm{d}\xi^{2}}=\frac{1}{2}\left(\frac{\mathrm{d}g_{0}}{\mathrm{d}\xi}\right)^{2}-2g_{0}^{3}-\xi g_{0}^{2}+\frac{1}{2}A_{0}-\frac{1}{8},$ (3.18)
where the constant has been determined by enforcing the far-field condition (3.6). Equation (3.18) is known as Painlevé 34 (P34), since an equivalent version was labelled XXXIV in section 14.33 of Ince [57].
Clearly, given a solution $F$ of (3.14)–(3.16), we can recover a solution of P34 by combining (3.9) and (3.17) to give $g_{0}=\mathrm{d}F/\mathrm{d}\xi-F^{2}-\xi/2$. Therefore all of our results for $F$ could be framed in terms of $g_{0}$. Note the one-to-one correspondence between solutions of PII and P34 was given in Ref. [43], for example. Equivalent links between PII and P34 can be established via the associate Hamiltonian, as described in section 32.6(iii) in the DLMF [38] (equation 32.6.12 is equivalent to our P34 in (3.18)). See section 2.4 of Clarkson [24] for further details.
@@ 3.4 Liouville–Green (WKB) analysis and Stokes phenomenon
For simplicity, we assume for the moment that the initial condition $u_{0}(x)$ has a single double pole in the upper half plane at $x=x_{0}$ (like (1.6) does). We wish to study (3.14) subject to (3.15). The solution we are after will satisfy this far-field condition in a broader sector that includes the negative imaginary direction. To determine this sector, and also to determine whether (3.15) is sufficient to specify our solution uniquely, we shall linearise about (3.15).
@@
which is related to the modified Bessel’s equation via a change of variables. The two possible linearly independent solutions, namely
$\xi^{1/2}I_{(1+24\alpha^{2})^{1/2}/3}({\textstyle\frac{2}{3}}\xi^{3/2})\quad\mbox{and}\quad\xi^{1/2}K_{(1+24\alpha^{2})^{1/2}/3}({\textstyle\frac{2}{3}}\xi^{3/2}),$
where $I_{\nu}(z)$ and $K_{\nu}(z)$ are modified Bessel functions of the first and second kind, have the far-field behaviour $J\sim\mathrm{const}\,\xi^{-1/4}\,\mathrm{e}^{\pm 2\xi^{3/2}/3}$. Therefore we have
@@
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty.$ (3.19)
This exercise demonstrates that (3.15) is acting as one boundary condition only and to specify the solution of (3.14)–(3.15) uniquely we need to fix $\sigma_{1}$ in (3.19), which is hidden beyond all orders of the algebraic expansion.
At this stage, we emphasise that (3.14) with (3.19) combine to give a one-parameter family of solutions of the type studied at length by Boutroux [16] and by others, including in Refs [44, 46, 58, 60, 61]. We note that we are able to write out the full algebraic series in (3.19) as
$F\sim\frac{\alpha}{\xi}\sum_{n=0}^{\infty}\frac{b_{n}}{\xi^{3n}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (3.20)
@@
$b_{n+1}=(3n+1)(3n+2)b_{n}-2\alpha^{2}\sum_{\begin{subarray}{c}i,j,k=0\\ i+j+k=n\end{subarray}}^{n}b_{i}b_{j}b_{k},\quad b_{0}=1.$ (3.21)
@@
$b_{n}\sim\frac{\sin\pi\alpha}{\alpha\pi^{3/2}}\left(\frac{3}{2}\right)^{2n+1/2}\Gamma\left(2n+{\textstyle\frac{1}{2}}\right)\quad\mbox{as}\quad n\rightarrow\infty$ (3.22)
(see also [44]). A key point is that for $\alpha\in\mathbb{N}$, the series (3.20) is convergent (see section 3.5 below), while otherwise it is divergent.
To proceed with our argument about the one-parameter family of solutions to (3.14) with (3.19), we explore broader far-field sectors in the $\xi$ plane, referring to figure 7. With $\xi=\rho\mathrm{e}^{\mathrm{i}\theta}$, then let the direction $\xi\rightarrow-\mathrm{i}\infty$ be $\theta=-\pi/2$. For any $\sigma_{1}\neq 0$, then we would have $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ decaying along the ray $\theta=-\pi/2$. Along a path $\rho=\,\mathrm{const}$ (with $\rho\gg 1$), as $\theta$ increases from $\theta=-\pi/2$ this term would increase in size until it was $\mathcal{O}(1)$ at the anti-Stokes line $\theta=-\pi/3$ (dashed line in figure 7). For $\theta>-\pi/3$, the term would be exponentially growing.
@@
In order to eliminate the inclusion of an exponentially growing term in the sector $\pi<\theta<0$, we need to force $\sigma_{1}=0$ in (3.19). In that case, both exponentials $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ and $\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$ are absent along $\theta=-\pi/2$. As $\theta$ increases from $\theta=-\pi/2$, we see that $\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$ is switched on at the Stokes line $\theta=0$ (solid blue line) and becomes of $\mathcal{O}(1)$ at $\theta=\pi/3$. Transversing the other way, as $\theta$ decreases from $\theta=-\pi/2$, the exponential $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ is switched on at the Stokes line $\theta=-2\pi/3$ (solid blue line) and becomes of $\mathcal{O}(1)$ at $\theta=-\pi$. Putting it together, a complete description of our inner problem is
$\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}=2F^{3}+\xi F-\alpha,$ (3.23)
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad|\xi|\rightarrow\infty,\quad-2\pi/3<\mathrm{arg}(\xi)<0$ (3.24)
$\sigma_{1}=0,\quad-2\pi/3<\mathrm{arg}(\xi)<0,$ (3.25)
where it is understood that the ellipsis represent further terms in the divergent power series (3.20); that is, there is a one-parameter family of solutions that satisfies (3.23)–(3.24), but the specific solution we are after is selected by enforcing (3.25). Furthermore, the unique solution to (3.23)–(3.25) should be pole-free region in the far field of the sector $\pi<\theta<\pi/3$ (bounded by active anti-Stokes lines at $\theta=-\pi$ and $\theta=\pi/3$) and the power-series part of the far-field condition, namely
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}\quad\mbox{as}\quad|\xi|\rightarrow\infty,$ (3.26)
@@
Finally, from a practical (numerical) perspective, perhaps the best way to enforce (3.24)–(3.25) (which count as two boundary conditions) is to apply the far-field condition (3.26) along the ray $\theta=-\pi/3$ (the inactive anti-Stokes line). The reason is that for any $\sigma_{1}\neq 0$, the condition (3.26) would not apply along $\theta=-\pi/3$ (since, as already explained, the term $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ would be $\mathcal{O}(1)$ along this ray). Therefore, applying (3.26) along $\theta=-\pi/3$ has the effect of killing off $\sigma_{1}\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ (as well as the other exponential $\sigma_{2}\xi^{-1/4}\mathrm{e}^{-2\xi^{3/2}/3}$). As an example of a numerical solution of (3.23)–(3.25) that is found by enforcing (3.26) along $\theta=-\pi/3$, we show in figure 8 an image computed for the case $\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{5})$ using the algorithm outlined by Fornberg & Weideman [46]. Here the yellow dots indicate poles of the solution, which are aligned in arrays. The red lines indicate anti-Stokes lines. Note that for this value of $\alpha$, there are no poles in the third or fourth quadrant. We discuss other numerical solutions of (3.23)–(3.25) below in subsection 3.6.
@@
Figure 8: An image showing $|F|$ for a numerical solution of Painlevé II with $\alpha={\textstyle\frac{1}{2}}(-1+\sqrt{5})$ using the algorithm from Fornberg & Weideman [46]. The yellow dots represent the pole field for this solution. The two active anti-Stokes lines are shown in red. While solutions of PII can be uniquely specified by the “initial conditions” $F(0)$ and $F^{\prime}(0)$, our scheme provides this pair as outputs. For this example, the numerically obtained values are $F(0)\approx 0.5941+1.0289\mathrm{i}$ and $F^{\prime}(0)\approx 0.7995-1.3848\mathrm{i}$.